is 3y+2x=0 parallel, perpendicular, or neither to y=1.5x
Answer by stanbon(75887)

is 3y+2x=0 parallel, perpendicular, or neither to y=1.5x
----
Look at the slopes.
-------
1st equation: slope = -2/3
2nd equation: slope = 1.5 = 3/2
--------
These lines are perpendicular because
the product of their slopes = -1.
====================================
Cheers,
Stan H.
===============